31 January 2010 Technical Night 2 --- continued LBC Active Optics testing & LUCI prep (20100131)

Wayne Wack, Dave Thompson at LBT Andrew Rakich, Olga Kuhn at Steward

06:30 About 11pm or 11:30, opened and pointing check done

07:20 Running dofpia - seeing ~1-1.5"

dofpia run --- bluelimit=400 (default). Looking at Z4, Z22, Z11 convergence.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1111    275  -1284     54    135    -27     95    y  1.10  0.59                      /newdata/lbcb.20100131.071219.fits  V-BESSEL
   586   -238   -391    -12     27   -346     21    y  1.04  0.35                      /newdata/lbcb.20100131.071345.fits  V-BESSEL
  1200    -18    562   -412    263   -341     14    y  1.05  0.59                      /newdata/lbcb.20100131.071504.fits  V-BESSEL
  -528      7   -391   -119   -147    -41     71    y  1.25  0.27                      /newdata/lbcb.20100131.071634.fits  V-BESSEL
   910   -163    344     32    -51   -341    -11    y  1.02  0.44                      /newdata/lbcb.20100131.071757.fits  V-BESSEL
   532   -219    741    159    464     -2     48    y  1.52  0.39                      /newdata/lbcb.20100131.071919.fits  V-BESSEL
   265   -198    620     -9    102    -17     17    y  1.90  0.24                      /newdata/lbcb.20100131.072048.fits  V-BESSEL
  1835      0      0      0      0   -966      0    y  0.00  0.00                     dofpia_sph3_adjustment             

Now with bluelimit=200 (but note the poor seeing estimate of 3.33" compromised this)

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -832     74   -399   -215   -149     71     18    y  1.29  0.38                      /newdata/lbcb.20100131.072312.fits  V-BESSEL
   222   -233   -567     39   -396      0      3    y  2.39  0.28                      /newdata/lbcb.20100131.072440.fits  V-BESSEL
   467      6   -373    273    253   -204    -36    y  1.23  0.31                      /newdata/lbcb.20100131.072608.fits  V-BESSEL
  -817    178    320    -33    -26    371      7    y  1.34  0.42                      /newdata/lbcb.20100131.072731.fits  V-BESSEL
   882   -240   -269     95     36   -173    -35    y  2.08  0.39                      /newdata/lbcb.20100131.072855.fits  V-BESSEL
   536   -354     96     65     -1   -122    -21    y  1.34  0.26                      /newdata/lbcb.20100131.073020.fits  V-BESSEL
   440    422    -69   -262     22      0      0    y  3.33  0.26                      /newdata/lbcb.20100131.073145.fits  V-BESSEL
  2782      0      0      0      0  -1464      0    y  0.00  0.00                     dofpia_sph3_adjustment             

07:30 We are starting a test of the off-axis SW, but it is not finding any tech chip pupils. I see some on the tech chip images #2 but these have large negative spherical, possibly because the last fpia image which 'converged' did so with a huge seeing value (3.33) and hence a large negative z11 was sent. We do need to plan these pointings better so bright stars will be on the AO tech chip.

07:45 RH was rising sharply and around 00:45 MST reached 95% so we had to close. They will reconfigure to LUCIFER during this downtime.

-- OlgaKuhn - 31 Jan 2010
Topic revision: r1 - 31 Jan 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback